OpenStudy (anonymous):

At midnight, the temperature was -12°F. By noon the temperature had risen 25°F. During the afternoon, if fell 10°F and fell another 3°F by midnight. What was the final temperature? is it 0?

OpenStudy (angela210793):

-12+25-10-3=____

OpenStudy (anonymous):

13-10=3-3=0

OpenStudy (angela210793):

:)

OpenStudy (anonymous):

0

OpenStudy (anonymous):

thanks